OpenStudy (anonymous):

At the Boston Children's Museum, it was estimated 38% of all visitors are from in state. On Saturday, 8,000 people attended the museum. What is the number of out of state people in attendance? (Points : 2)

OpenStudy (anonymous):

it is 62% people out of state ... so 800*6 +800/5 = 4800+160 = 4960

OpenStudy (anonymous):

All you have to do is take 8,000 * .38, and then subtract that number from 8,000. Which you will get 4960.